Q5. A column foundation shown below is 3mx2m in plan. Given: D=2m, =20°, and c=45 kN/m2. Using Meyerhofs bearing capacity

Please solve step by step and clearly thank you
Answer #1

N=16 kN/m Groundwateę 2M & sat = 19.4 kN/ 3 mx 2m O= 20 c = u5kN/m² FS=3 According to Meyerhoffs general beering Capacity e14.83 fes, fqs, Frs > Shape factors fcs = 1+ B Nq = 1+ ? x 6.4 Fes = 1.288 fas= 1+ B Tan q = 1+ 2 Tanz . fqs = 1.243 Frs = 1-1.187 Fed = frd = q= overbueden stress = roux Df 9= 16.8x1 + 1*(19.4-9.01) q= 26.39 kN/m² r= Submerged unit weight = Vsat - Nu = ( 3 2 8 IN / Net ultimate bearing capacity - quet = qu - Doux Df = 1301.28 - ( 16.8x1 + (19.4-9.01)x)) - quet = 1274.89 k
Answer #2

wrong, recheck your formula